Find the percentile rank of 41.
Data Set : 23, 54, 32, 34, 45, 41, 47, 38, 28, 71, 36, 19, 26, 31
Answer #1

Solution:
Arranging Observations in the ascending order, We get :
19,23,26,28,31,32,34,36,38,41,45,47,54,71

Here, n=14

P41=(41(n+1)100)th value of the observation

=(41⋅15/100)th value of the observation

=(6.15)th value of the observation

=6th observation +0.15[7th-6th]

=32+0.15[34-32]

=32+0.15(2)

=32+0.3

=32.3